# Advancing Graph Generation through Beta Diffusion

### v1

@@ Abstract
Diffusion models have demonstrated effectiveness in generating natural images and have been extended to generate diverse data types, including graphs. This new generation of diffusion-based graph generative models has demonstrated significant performance improvements over methods that rely on variational autoencoders or generative adversarial networks. It’s important to recognize, however, that most of these models employ Gaussian or categorical diffusion processes, which can struggle with sparse and long-tailed data distributions. In our work, we introduce Graph Beta Diffusion (GBD), a diffusion-based generative model particularly adept at capturing diverse graph structures. GBD utilizes a beta diffusion process, tailored for the sparse and range-bounded characteristics of graph adjacency matrices. Furthermore, we have developed a modulation technique that enhances the realism of the generated graphs by stabilizing the generation of critical graph structures, while preserving flexibility elsewhere. The outstanding performance of GBD across three general graph benchmarks and two biochemical graph benchmarks highlights its capability to effectively capture the complexities of real-world graph data. The code will be made available on GitHub. The code will be made available at https://github.com/YH-UtMSB/Graph_Beta_Diffusion.
@@ 1 Introduction
In recent years, the field of machine learning-driven graph generation has witnessed a surge in interest and activity. This growing attention is driven by the recognition of graph data’s pervasive presence and utility across diverse real-world applications, ranging from social network study (Newman et al., 2002; Conti et al., 2011; Abbe, 2018; Shirzadian et al., 2023) to biochemical molecular research (Jin et al., 2018; Liu et al., 2018; Bongini et al., 2021; Guo et al., 2024; Li and Yamanishi, 2024). Additionally, the rapid evolution of machine learning tools has introduced powerful techniques for data generation, among which diffusion models (Ho et al., 2020; Song et al., 2021; Austin et al., 2021; Avdeyev et al., 2023; Chen and Zhou, 2023; Zhou et al., 2023) stand out as a notable example. As these advanced tools intersect with the task of graph generation, we witness the emergence of numerous diffusion-based graph generative models (Niu et al., 2020a; Jo et al., 2022; Haefeli et al., 2022; Huang et al., 2022; Vignac et al., 2023; Jo et al., 2023; Cho et al., 2023; Chen et al., 2023; Kong et al., 2023).
@@
While diffusion-based graph generative models often demonstrate superior performance compared to their predecessors (You et al., 2018; De Cao and Kipf, 2018; Li et al., 2018; Simonovsky and Komodakis, 2018; Liu et al., 2019; Shi et al., 2020; Luo et al., 2021; Martinkus et al., 2022), there is still potential for further enhancement in the quality of generated graphs. Among the latest advancements in these methods, it is widely recognized that incorporating inductive bias from the graph data is generally beneficial for model design (Jo et al., 2023). One promising direction of incorporating this bias involves considering the statistical characteristics of the distribution of graph data. For instance, both Graph D3PM (Haefeli et al., 2022) and DiGress (Vignac et al., 2023) have demonstrated that when considering the binary or categorical nature of the graph adjacency matrix and modeling it in the discrete space, it provides benefits for generating more realistic graphs.
@@
Accounting for the discreteness of the graph adjacency matrix has shown enhancement to model performance. However, it is crucial to recognize that the complexity and flexibility of distribution characteristics of graph data extend beyond mere discreteness. Real-world graphs typically exhibit sparsity in edge distribution and long-tailedness in nodal statistical properties such as degrees and categories (Barabási et al., 2000; Ciotti et al., 2015; Liang et al., 2023; Wang et al., 2023). Moreover, when mapping the graph structure into an adjacency matrix, the values within the matrix are also bounded by the range of edge weights. Given these unique characteristics inherent to graph data, it becomes apparent that Gaussian and categorical distributions, which are typically the default choices for building the diffusion processes, may not align well with these graph characteristics. Consequently, this mismatch could potentially introduce limitations when modeling the distribution over graphs.
@@
Considering the desirable statistical characteristics of graph data, we find that the beta distribution emerges as a particularly suitable modeling choice. With properties such as being range-bounded and flexible to model data at all sparsity levels, the beta distribution aligns well with the inherent traits of graphs, hence making itself a promising candidate to surpass the potential limitations imposed by utilizing Gaussian or categorical distributions. In this paper, we introduce Graph Beta Diffusion (GBD) as a novel addition to diffusion-based graph generative models. GBD models the joint distribution of node attributes and edge connections within a graph through beta diffusion (Zhou et al., 2023), a generative diffusion process developed upon the thinning and thickening of beta random variables. In the forward diffusion process, the original data is gradually attenuated toward zero. This process starts with a modified form of the data, which is scaled and shifted to ensure it is bounded between 0 and 1. At each step, a random proportion, sampled from a beta distribution with a predefined noise schedule, is multiplied with the attenuated data. Conversely, the reverse diffusion process starts from zero and attempts to recover the original data. It does this by adding a random proportion of the difference between one and the current value to the current value at each step.
@@ 2 The Methodology
Figure 1: Overview of GBD. The beta random variables $Q_{t}$ and $P_{t}$ are sampled from the beta distribution parameterized by $G_{0}$ and clean graph predicted by the neural network $\tilde{G}_{\theta}$, respectively. $\tilde{G}_{\theta}$ is learned through minimizing $\mathcal{L}_{\mathrm{sampling}}$ and $\mathcal{L}_{\mathrm{correction}}$ defined in Section 2.3.
2.1 Data description and mathematical notations
In this study, our primary focus lies in generating two types of graphs: generic graphs and molecular graphs. Generic graphs are characterized as undirected, simple graphs with $N$ nodes. Each pair of nodes $(u,v)$ in these graphs can only have two possible edge statuses: connected or disconnected. Therefore, the entire graph structure can be fully described by a symmetric binary adjacency matrix ${\mathbf{A}}\in\{0,1\}^{n\times n}$, where “$1$”s denote connected node pairs and “00”s denote pairs that are not directly connected. Molecular graphs are also simple graphs, but they typically feature multiple types of edges. As a result, their graph structures need to be represented as adjacency matrices with dummy-encoded categorical variable elements, which is a common practice established in previous research (Jo et al., 2022; Vignac et al., 2023; Jo et al., 2023). This expression results in a multitude of channels in the graph adjacency matrix, we use ${\mathbf{A}}^{(1:K)}$ to represent the structure of a graph with $K$ types of edges.
While the primary goal of graph generation and quality evaluation centers around graph architecture, incorporating the node features can be advantageous for learning the distribution of the adjacency matrices. Node features are typically represented by a matrix ${\mathbf{X}}$ of shape $n\times d$, where $d$ denotes the number of features. The choices of node features offers high flexibility, ranging from raw-data-provided node categories to some hand-crafted features, such as node-level statistics (Jo et al., 2022) or spectral graph signals (Jo et al., 2023). The features within ${\mathbf{X}}$ exhibit great diversity in their nature, including numerical, categorical, and ordinal types. Through preprocessing methods including dummy-encoding and empirical CDF transformation, we standardize them as continuous variables bounded by $[0,1]$. In this section, we denote the target datum as ${\mathbf{G}}$. For generic graphs, ${\mathbf{G}}$ comprises $({\mathbf{A}},{\mathbf{X}})$; for molecular graphs, ${\mathbf{G}}$ is defined as $({\mathbf{A}}^{(1:K)},{\mathbf{X}})$. In the sequel, We by default employ the generic graph scenario to illustrate the methodology.
@@ 2.2 Forward and reverse beta diffusion processes
Forward beta diffusion process.
Such a process can be characterized by the transition probability $q({\mathbf{G}}_{t}|{\mathbf{G}}_{t-1},{\mathbf{G}}_{0})$, with ${\mathbf{G}}_{0}$ denoting the combination of the original adjacency matrix and node feature matrix. Following recent diffusion-based graph generative models (Jo et al., 2022; Vignac et al., 2023; Jo et al., 2023; Cho et al., 2023), we assume $q({\mathbf{G}}_{t}|{\mathbf{G}}_{t-1},{\mathbf{G}}_{0})$ to be factorizable such that $q({\mathbf{G}}_{t}|{\mathbf{G}}_{t-1},{\mathbf{G}}_{0})=q({\mathbf{A}}_{t}|{\mathbf{A}}_{t-1},{\mathbf{A}}_{0})\cdot q({\mathbf{X}}_{t}|{\mathbf{X}}_{t-1},{\mathbf{X}}_{0})$. Constructing the forward beta diffusion process (Zhou et al., 2023) for graph modeling, we have:
@@
Reverse beta diffusion process.
It is important to note that the joint distribution $q({\mathbf{G}}_{1:T}|{\mathbf{G}}_{0})$ can be equivalently constructed in reverse order through ancestral sampling, which directs samples from the terminus states ${\mathbf{G}}_{T}$ towards the initial states ${\mathbf{G}}_{0}$ by incrementally applying the changes $\delta{\mathbf{G}}_{t}$ at each reversed time stamp. With the changes at a given time $t$ parameterized as $\delta{\mathbf{G}}_{t}:={\mathbf{P}}_{t}\odot(1-{\mathbf{G}}_{t})$, where ${\mathbf{P}}_{t}$ are beta-distributed fractional multipliers, the time-reversal sampling process can be mathematically defined as: for $t=T,T-1,\cdots,1$,

### v2

@@ Abstract
Diffusion models have excelled in generating natural images and are now being adapted to a variety of data types, including graphs. However, conventional models often rely on Gaussian or categorical diffusion processes, which can struggle to accommodate the mixed discrete and continuous components characteristic of graph data. Graphs typically feature discrete structures and continuous node attributes that often exhibit rich statistical patterns, including sparsity, bounded ranges, skewed distributions, and long-tailed behavior. To address these challenges, we introduce Graph Beta Diffusion (GBD), a generative model specifically designed to handle the diverse nature of graph data. GBD leverages a beta diffusion process, effectively modeling both continuous and discrete elements. Additionally, we propose a modulation technique that enhances the realism of generated graphs by stabilizing critical graph topology while maintaining flexibility for other components. GBD competes strongly with existing models across multiple general and biochemical graph benchmarks, showcasing its ability to capture the intricate balance between discrete and continuous features inherent in real-world graph data. The PyTorch code is available on GitHub.
Figure 1: Overview of the forward and reverse diffusion processes of GBD. The multiplicative factors ${\mathbf{Q}}_{t}$ and ${\mathbf{P}}_{t}$ are sampled from beta distributions parameterized by the initial graphs ${\mathbf{G}}_{0}$ and “clean graphs” predicted by $\hat{G}_{\theta}$. The neural network $\hat{G}_{\theta}$ is learned through minimizing Equation 8 constituted by $\mathcal{L}_{\mathrm{sampling}}$ and $\mathcal{L}_{\mathrm{correction}}$.
Figure 2: Edge generation process of the GBD model in graph topology (top) and adjacency matrix (bottom) views. Nodes are sorted by descending degree centrality and color-coded by degree, from yellow (high) to purple (low). The modulation technique in Section 2.3 enables early emergence of key positions such as community hubs, enhancing reverse chain stability.
@@ 1 Introduction
In recent years, there has been a significant surge in interest and activity in the field of graph generation, particularly with the development of advanced generative models tailored for graphs. This growing attention is driven by the recognition of graph data’s pervasive presence and utility across diverse real-world applications, ranging from social network study (Newman et al., 2002; Conti et al., 2011; Abbe, 2018; Shirzadian et al., 2023) to biochemical molecular research (Jin et al., 2018; Liu et al., 2018; Bongini et al., 2021; Guo et al., 2024; Li & Yamanishi, 2024). Additionally, the rapid evolution of machine learning tools has introduced powerful techniques for data generation, among which diffusion models (Ho et al., 2020; Song et al., 2021; Austin et al., 2021; Avdeyev et al., 2023; Chen & Zhou, 2023; Zhou et al., 2023) stand out as a notable example. As these advanced tools intersect with the task of graph generation, we witness the emergence of numerous diffusion-based graph generative models (Niu et al., 2020a; Jo et al., 2022; Haefeli et al., 2022; Huang et al., 2022; Vignac et al., 2023; Jo et al., 2023; Cho et al., 2023; Chen et al., 2023; Kong et al., 2023).
@@
While diffusion-based graph generative models often demonstrate superior performance compared to their predecessors (You et al., 2018; De Cao & Kipf, 2018; Li et al., 2018; Simonovsky & Komodakis, 2018; Liu et al., 2019; Shi et al., 2020; Luo et al., 2021; Martinkus et al., 2022), there is still potential for further enhancement in the quality of generated graphs. Among the latest advancements in these methods, it is widely recognized that incorporating inductive bias from the graph data is generally beneficial for model design (Jo et al., 2023). One promising direction of incorporating this bias involves considering the statistical characteristics of the distribution of graph data. For instance, both Graph D3PM (Haefeli et al., 2022) and DiGress (Vignac et al., 2023) have demonstrated that when considering the binary or categorical nature of the graph adjacency matrix and modeling it in the discrete space, it provides benefits for generating more realistic graphs.
@@
Accounting for the discreteness of the graph adjacency matrix has shown enhancement to performance. However, it is crucial to recognize that the complexity and flexibility of the distribution characteristics of graph data extend beyond mere discreteness. Real-world graphs usually display sparse edge distributions and exhibit diverse statistical patterns in node attributes, which may include skewed, multi-modal, or long-tailed distributions (Barabási et al., 2000; Ciotti et al., 2015; Liang et al., 2023; Wang et al., 2023). While the values within node feature matrices may not inherently bounded by range, they are often be empirically represented or processed into quantities that are bound by specific limits. Considering the unique characteristics inherent to graph data, it is clear that Gaussian and categorical distributions, often default choices for constructing diffusion processes, may not adequately align with these graph traits. This misalignment could introduce noticeable limitations in accurately modeling the distribution of graphs.
@@
Given the unique statistical characteristics of graph data, the beta distribution emerges as a particularly suitable modeling choice. With great flexibility to model continuous data with various statistical characteristics and approximate discrete distributions at all sparsity levels, the beta distribution aligns well with the inherent traits of graphs, hence making itself a promising candidate to surpass the potential limitations imposed by utilizing Gaussian or categorical distributions. In this paper, we introduce Graph Beta Diffusion (GBD) as a novel addition to diffusion-based graph generative models. GBD models the joint distribution of node attributes and edge connections within a graph through beta diffusion (Zhou et al., 2023), a generative diffusion process that is built upon the thinning of beta random variables in its multiplicative forward diffusion process and the thickening in its multiplicative reverse process.
@@ 2 The Methodology
In this study, our primary focus lies in generating two types of graphs: generic graphs and molecular graphs. A graph with $N$ nodes is represented by the tuple ${\mathbf{G}}=({\mathbf{A}},{\mathbf{X}})$, where ${\mathbf{X}}\in\mathbb{R}^{N\times D}$ denotes the node features with feature dimension $D$, and ${\mathbf{A}}\in\mathbb{R}^{N\times N}$ is the symmetric binary adjacency matrix that defines the connections between nodes. The selection of node features offers high flexibility, ranging from raw-data-provided node categories to hand-crafted features such as node-level statistics (Jo et al., 2022) or spectral graph signals (Jo et al., 2023). The features within ${\mathbf{X}}$ exhibit great diversity in their nature, including numerical, categorical, and ordinal types. Through preprocessing methods including dummy-encoding, empirical CDF transformation or normalization, we standardize them as continuous variables bounded by $[0,1]$. For molecular graphs, we use ${\mathbf{A}}^{(1:K)}$ to represent the structure of a graph with $K$ types of edges and ${\mathbf{G}}$ is defined as $({\mathbf{A}}^{(1:K)},{\mathbf{X}})$. In the sequel, we by default employ the generic graph scenario to illustrate the methodology.
@@ 2.1 Forward and reverse beta diffusion processes
Forward multiplicative beta diffusion process.
Such a process can be characterized by the transition probability $q({\mathbf{G}}_{t}\,|\,{\mathbf{G}}_{t-1},{\mathbf{G}}_{0})$, with ${\mathbf{G}}_{0}$ denoting the combination of the original adjacency matrix and node feature matrix. Following recent diffusion-based graph generative models (Jo et al., 2022; Vignac et al., 2023; Jo et al., 2023; Cho et al., 2023), we assume $q({\mathbf{G}}_{t}\,|\,{\mathbf{G}}_{t-1},{\mathbf{G}}_{0})$ to be factorizable such that $q({\mathbf{G}}_{t}\,|\,{\mathbf{G}}_{t-1},{\mathbf{G}}_{0})=q({\mathbf{A}}_{t}\,|\,{\mathbf{A}}_{t-1},{\mathbf{A}}_{0})\cdot q({\mathbf{X}}_{t}\,|\,{\mathbf{X}}_{t-1},{\mathbf{X}}_{0})$. Constructing the forward multiplicative beta diffusion process (Zhou et al., 2023) for graph modeling, we have:
@@
Reverse multiplicative beta diffusion process.
It is important to note that the joint distribution $q({\mathbf{G}}_{1:T}\,|\,{\mathbf{G}}_{0})$ can be equivalently constructed in reverse order through ancestral sampling, which directs samples from the terminus states ${\mathbf{G}}_{T}$ towards the initial states ${\mathbf{G}}_{0}$ by incrementally applying the changes $\delta{\mathbf{G}}_{t}$ at each reversed time stamp. With the changes at a given time $t$ parameterized as $\delta{\mathbf{G}}_{t}:={\mathbf{P}}_{t}\odot(1-{\mathbf{G}}_{t})$, where ${\mathbf{P}}_{t}$ are beta-distributed fractional multipliers, the time-reversal multiplicative sampling process can be mathematically defined as: for $t=T,T-1,\cdots,1$,